40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Aug 1 01:13:32 2008, rana, Configuration, PSL, PSL Quad change and new script 
    Reply  Fri Aug 1 16:33:52 2008, rana, Configuration, PSL, PSL Quad change and new script Untitled.png
Message ID: 778     Entry time: Fri Aug 1 01:13:32 2008     Reply to this: 781
Author: rana 
Type: Configuration 
Category: PSL 
Subject: PSL Quad change and new script 
Koji and I changed a few optics so that now ~60% of the beam that went to the PSL POS QPD
now goes to the west side of the table for the aux. laser locking PLL. The beam is sort of
on the QPD again but needs a centering.

After this work I wrote a script SUS/freeswing-all.csh which puts a 30000 count offset into
the UL coil of each suspension and then disables it. This is just good enough to kick it up
so that the eigenfrequency can be measured. I ran it and it worked -- it finished running at

Fri Aug 1 00:44:30 PDT 2008
ELOG V3.1.3-